OpenStudy (anonymous):

Integral of (3x-2)/(x-1)^2

OpenStudy (anonymous):

Using partial fractions Idk how to successfully get my A value

myininaya (myininaya):

\[\frac{3x-2}{(x-1)^2}=\frac{A}{x-1}+\frac{Bx}{(x-1)^2}=\frac{A(x-1)+Bx}{(x-1)^2}\] 3x-2=A(x-1)+Bx 3x-2=x(A+B)-A A+B=3 -A=-2=>A=2 A+B=3=>2+B=3=>B=1 lets check our partial fraction we have \[\frac{2}{x-1}+\frac{x}{(x-1)^2}=\frac{2(x-1)+x}{(x-1)^2}=\frac{3x-2}{(x-1)^2}\] so yay no mistakes!

OpenStudy (anonymous):

Why is it Bx and not just A + B

OpenStudy (amistre64):

it can be just B since they are mulitples of linears

OpenStudy (amistre64):

but it works out anyways right?

OpenStudy (anonymous):

Do you mean it can or can't?

OpenStudy (amistre64):

when in doubt; try it both ways ... unless its illegal :)

OpenStudy (anonymous):

And it is just B.

myininaya (myininaya):

\[\int\limits_{}^{}\frac{3x-2}{(x-1)^2}dx=\int\limits_{}^{}(\frac{2}{x-1}+\frac{x}{(x-1)^2})dx=2\ln|x-1|+\int\limits_{}^{}\frac{x}{(x-1)^2}dx\] let u=x-1 so du=1dx also if u=x-1, then x=u+1 so we have \[2\ln|x-1|+\int\limits_{}^{}\frac{u+1}{u^2}du=2\ln|x-1|+\int\limits_{}^{}(\frac{1}{u}+\frac{1}{u^2})du=2\ln|x-1|+\ln|u|-\frac{1}{u}\+C] =\[2\ln|x-1|+\ln|x-1|-\frac{1}{x-1}+C=\ln|x-1|^2+\ln|x-1|-\frac{1}{x-1}+C\] =\[\ln|{(x-1)^2(x-1)}|-\frac{1}{x-1}+C=\ln|(x-1)^3|-\frac{1}{x-1}+C\] you can use this to check yourself after you do the problem if you like Check: \[(\ln|(x-1)^3|-\frac{1}{x-1}+C)'=3*\frac{1}{x-1}+\frac{1}{(x-1)^2}+0=\frac{3(x-1)+1}{(x-1)^2}=\frac{3x-1+1}{(x-1)^2} =\frac{3x}{(x-1)^2}\] so i made a mistake somewhere

myininaya (myininaya):

my thingy got cut off